# Two-Player Zero-Sum Games with Bandit Feedback

### v2

@@ Appendix C Proof of Theorem 2
According to the algorithm, $4\sigma\sqrt{e/T}$ is a critical value for $\hat{\Delta}_{t}$ since it takes this value in round $\frac{1}{2}\log_{2}\frac{T}{e}$. In this way, let define a threshold parameter as $\lambda\geq 4\sigma\sqrt{e/T}$, which gives us two subsets of actions pairs such that $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$. Since $\Delta_{ij}^{*}\leq\Delta_{ij}$ ensures the bounds are preserved, it makes sense to use these action pair sets in the analysis. Thus, we can write the Nash regret as
$R_{T}^{*}=\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\mathbb{E}[n_{ij,T}]+\sum_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}\mathbb{E}[n_{ij,T}]$
where $\Delta_{ij}^{*}=A(i^{*},j^{*})-A(i,j).$ and $n_{ij,T}$ is the total number of times to play action pair $(i,j)$ by time $T$.
To analyze the Nash regret of Algorithm 2, it is essential to consider various scenarios that may lead to suboptimal outcomes. Each of these scenarios contributes to the total Nash regret, as they either involve the incorrect elimination of the optimal action pair or the unnecessary selection of suboptimal pairs. The former results in discarding the optimal strategy because of insufficient evidence, while the latter leads to spend more time with action pairs that are unlikely to be part of any near optimal equilibrium.
The algorithm employs a scheduling mechanism for $\Delta_{ij}$, where $\hat{\Delta}_{t}$ is halved in each round. This approach is necessary because the true values of $\Delta_{ij}$ are unknown to the players. We use $\hat{\Delta}_{t}$ to determine both the exploration time and the tolerance level in the $\varepsilon$-NE property. Specifically, when $\hat{\Delta}_{t}$ is large, the algorithm plays each action pair fewer times and keeps more pairs during the elimination step, since the threshold for elimination is relatively loose. As $\hat{\Delta}_{t}$ decreases, the algorithm allocates more exploration to action pairs that are closer to being part of a NE, because those that are clearly suboptimal have already been eliminated. This adaptive process enables a more precise identification of near-optimal strategies by $\varepsilon$-NE condition while minimizing regret.
For each suboptimal action pair $(i,j)$, let $t_{ij}=\min\{t:\hat{\Delta}_{t}<\Delta_{ij}/2\}$ refer to the earliest round such that $\hat{\Delta}_{t}<\Delta_{ij}/2$. Using the fact that $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ and the definition of $t_{ij}$, we can write the following:
@@
Since neither some suboptimal nor the optimal action pair is not eliminated, the algorithm fails to discard suboptimal choices. It might converge to an incorrect solution without eliminating suboptimal actions, but since it still maintains the optimal action pair within the action pair selection set. The regret contribution of this case comes from the fact that the algorithm spends time with a worse choice when a better one is already available.
Let explain briefly in which situations the algorithm eliminates or keeps an action pair. For action pair $(i,j)$, if $\varepsilon$-NE property does not hold in round $t=t_{ij}$ which implies that at least one of the inequalities fails, then it will be eliminated in round $t_{ij}$.
Furthermore, we have $\varepsilon_{t_{ij}}=\sqrt{\frac{4\sigma^{2}}{k_{t_{ij}}}\ln{\Big(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}}\Big)}}\leq\frac{\hat{\Delta}_{t_{ij}}}{2}=\hat{\Delta}_{t_{ij}+1}<\frac{\Delta_{ij}}{4}$. In other words, if there exists $i^{\prime}\in S_{x}$ or $j^{\prime}\in S_{y}$, such that one of the following is true:
$I_{1}:\hat{A}(i^{\prime},j)-\hat{A}(i,j)>\varepsilon_{t_{ij}},\qquad I_{2}:\hat{A}(i,j)-\hat{A}(i,j^{\prime})>\varepsilon_{t_{ij}},$
then $(i,j)$ is eliminated. The inequality $I_{1}$ indicates that there exists an alternative action $i^{\prime}$ for the row player that offers a higher payoff than action $i$ against the column action $j$. Similarly, $I_{2}$ implies that there exists a better action $j^{\prime}$ for the column player than action $j$ when playing against the row action $i$. If either $I_{1}$ or $I_{2}$ holds, the action pair $(i,j)$ cannot be a part of the NE, as at least one player has an incentive to deviate. Hence, $(i,j)$ is not a NE with high probability.
@@
$\mathbb{P}(E_{1})$ $=\mathbb{P}(\hat{A}(i^{\prime},j)-\hat{A}(i,j)\leq\varepsilon_{t},\forall i^{\prime}\in S_{x})$ (52)
$\leq\mathbb{P}(\hat{A}(i^{\prime},j)-\hat{A}(i,j)\leq\varepsilon_{t})$ (53)
$\leq\exp{\Bigg(-\frac{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{\frac{4\sigma^{2}}{k_{t}}}\Bigg)}$ (54)
$=\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}$ (55)
where $\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}.$ To derive inequality (53), we use the fact that the probability of the intersection of multiple events is always less than or equal to the probability of any of individual events. Specifically, since we have $m$ actions in $S_{x}$, for any collection of events $e_{1},e_{2},\cdots,e_{m}$, it holds that $\mathbb{P}(e_{1}\cap e_{2}\cap\cdots\cap e_{m})\leq\mathbb{P}(e_{i})$ for all $i=1,2,\cdots,m$. In the last inequality (54), we apply Theorem 5 with $\sqrt{2\sigma^{2}/k_{t}}$-subgaussian random variable. The subgaussian parameter is derived by using Lemma 1 and we note that each action pair is played $k_{t}$ times.
@@
$\mathbb{P}(E_{2})=\mathbb{P}(\hat{A}(i,j)-\hat{A}(i,j^{\prime})\leq\varepsilon_{t},\forall j^{\prime}\in S_{y})\leq\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}.$ (56)
Therefore, we have the probability of keeping an action pair $(i,j)$ in any round $t$ as
$\mathbb{P}(E_{1}\cap E_{2})\leq\min\{\mathbb{P}(E_{1}),\mathbb{P}(E_{2})\}=\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}.$ (57)
It means that the probability that a suboptimal action pair $(i,j)$ is not eliminated in round $t_{ij}$ or before is bounded by $\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}T}$. Then, the regret contribution is simply bounded by a worst case which is $T\Delta_{ij}^{*}$ for any suboptimal action pair $(i,j)\in S_{A_{1}}$.
@@
$k_{t_{ij}}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil\leq\Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil.$
@@
It leads to a misidentification of the NE, implying that an action pair $(i,j)$ should be better than the optimal one based on the current estimates. By $\varepsilon$-NE property, clearly we have
@@
However, it is not enough to keep a suboptimal action pair $(i,j)$. On the other hand, we note that if the optimal action pair $(i^{*},j^{*})$ is eliminated in round $t_{*}$, it must fail to satisfy $\varepsilon$-NE property as the following:
@@
The optimal action pair can only be eliminated by a suboptimal action pair $(i,j)$ in round $t_{*}$ such that $t_{ij}\geq t_{*}$ since action pair $(i,j)$ must remain under consideration at the elimination round of the optimal action pair. This condition also implies that $\hat{\Delta}_{t_{ij}}\leq\hat{\Delta}_{t_{*}}$ because of the decreasing behavior of $\hat{\Delta}_{t}$ across rounds in the algorithm. Moreover, it is important to note that the algorithm keeps this action pair $(i,j)$ for at least one additional round following the elimination of the optimal pair as it is identified as a NE. We further observe that all action pairs $(i,j)$ with $t_{ij}<t_{*}$ have already been eliminated in round $t_{ij}$ or earlier, which we already consider this condition in the previous case.
@@
If the algorithm keeps a suboptimal action pair $(i,j)$ up to round $t_{*}$, the events specified in (51) must hold with parameter $\varepsilon_{t_{*}}$. Consequently, by applying (57), the probability of keeping a suboptimal action pair can be bounded by $\frac{16\sigma^{2}}{\hat{\Delta}_{t_{*}}^{2}T}$. Thus, we can write the regret contribution of this case as
$\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}T\frac{16\sigma^{2}}{\hat{\Delta}_{t_{*}}^{2}T}$ $<\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}T\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}+1}^{2}T}$ (62)
$=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}T\frac{64\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}T}$ (63)
@@
where $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ and we apply (50). We use $\hat{\Delta}_{t+1}$ because the optimal action pair might be eliminated at some round $t_{*}$ such that $t_{*}\in[0,\max_{ij}t_{ij}]$. This implies that the optimal action pair can be eliminated by any $(i,j)$ no later than the last round that $(i,j)$ remains in the game. However, we must keep $(i,j)$ at least until the next round, since it is not removed in round $t_{*}$ and is selected as a misidentified NE.
@@
This implies that a suboptimal action pair is played during the committing phase of Algorithm 2 up to step $T$. Thus, we need to account for an additional regret contribution term given by
$\sum_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T.$ (65)

### v3

@@ Appendix C Proof of Theorem 2
According to the algorithm, $4\sigma\sqrt{e/T}$ is a critical value for $\hat{\Delta}_{t}$ since it takes this value in the round $\frac{1}{2}\log_{2}\frac{T}{e}$. In this way, let define a threshold parameter as $\lambda\geq 4\sigma\sqrt{e/T}$, which gives us two subsets of actions pairs such that $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$. Since $\Delta_{ij}^{*}\leq\Delta_{ij}$, it ensures that the bounds are preserved and it makes sense to use these action pair sets in the analysis. Thus, we can write the Nash regret as
$R_{T}^{*}=\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\mathbb{E}[n_{ij,T}]+\sum_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}\mathbb{E}[n_{ij,T}]$ (52)
where $\Delta_{ij}^{*}=A(i^{*},j^{*})-A(i,j).$ and $n_{ij,T}$ is the total number of times to play the action pair $(i,j)$ by time $T$.
To analyze the Nash regret of Algorithm 2, it is essential to consider various scenarios that may lead to suboptimal outcomes. Each of these scenarios contributes to the total Nash regret, as they either involve the incorrect elimination of the optimal action pair or the unnecessary selection of suboptimal pairs. The former results in discarding the optimal strategy because of insufficient evidence, while the latter leads to spending more time with action pairs that are unlikely to be part of any near optimal equilibrium.
The algorithm employs a scheduling mechanism for $\Delta_{ij}$, where $\hat{\Delta}_{t}$ is halved in each round. This approach is necessary because the true values of $\Delta_{ij}$ are unknown to the players. We use $\hat{\Delta}_{t}$ to determine both the exploration time and the tolerance level in the $\varepsilon$-NE property. Specifically, when $\hat{\Delta}_{t}$ is large, the algorithm plays each action pair fewer times and keeps more pairs during the elimination step, since the threshold for elimination is relatively loose. As $\hat{\Delta}_{t}$ decreases, the algorithm allocates more exploration to action pairs that are closer to being part of a NE, because those that are clearly suboptimal have already been eliminated. This adaptive process enables a more precise identification of near-optimal strategies by the $\varepsilon$-NE condition while minimizing regret.
For each suboptimal action pair $(i,j)$, let $t_{ij}=\min\{t:\hat{\Delta}_{t}<\Delta_{ij}/2\}$ refer to the earliest round such that $\hat{\Delta}_{t}<\Delta_{ij}/2$. Specifically, we are interested in the first round where $\hat{\Delta}_{t}$ satisfies both $\hat{\Delta}_{t}<\Delta_{ij}^{\text{max}}$ and $\hat{\Delta}_{t}<\Delta_{ij}^{\text{min}}$, but combining these conditions results in the simplified condition $\hat{\Delta}_{t}<\Delta_{ij}/2$. Using the fact that $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ and the definition of $t_{ij}$, we can write the following:
@@
Furthermore, we have the following:
$\varepsilon_{t_{ij}}=\sqrt{\frac{4\sigma^{2}}{k_{t_{ij}}}\ln{\Big(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}}\Big)}}\leq\frac{\hat{\Delta}_{t_{ij}}}{2}=\hat{\Delta}_{{t_{ij}}+1}<\frac{\Delta_{ij}}{4}$ (54)
where the exploration time $k_{t}$ is defined in (21), which satisfies $k_{t_{ij}}\geq\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}$.
@@
Since neither some suboptimal nor the optimal action pair is eliminated, the algorithm fails to discard suboptimal choices. It might converge to an incorrect solution without eliminating suboptimal actions, but since it still maintains the optimal action pair within the action pair selection set. The regret contribution of this case comes from the fact that the algorithm spends time with a worse choice when a better one is already available.
Let explain briefly in which situations the algorithm eliminates or keeps an action pair. For action pair $(i,j)$, if the $\varepsilon$-NE property does not hold in round $t=t_{ij}$, which implies that at least one of the inequalities fails, then it will be eliminated in round $t_{ij}$. In other words, if there exists $i^{\prime}\in S_{x}$ or $j^{\prime}\in S_{y}$, such that one of the following is true:
$I_{1}:\hat{A}(i^{\prime},j)-\hat{A}(i,j)>\varepsilon_{t},\qquad I_{2}:\hat{A}(i,j)-\hat{A}(i,j^{\prime})>\varepsilon_{t},$ (55)
then $(i,j)$ is eliminated. The inequality $I_{1}$ indicates that there exists an alternative action $i^{\prime}$ for the row player that offers a higher payoff than action $i$ against column action $j$. Similarly, $I_{2}$ implies that there exists a better action $j^{\prime}$ for the column player than the action $j$ when playing against the row action $i$. If $I_{1}$ or $I_{2}$ holds, the action pair $(i,j)$ cannot be part of the NE, as at least one player has an incentive to deviate. Hence, $(i,j)$ is not a NE with high probability.
@@
$\mathbb{P}(E_{1})$ $=\mathbb{P}(\hat{A}(i,j)-\hat{A}(i^{\prime},j)+\varepsilon_{t}\geq 0,\forall i^{\prime}\in S_{x})$ (57)
$\leq\mathbb{P}(\hat{A}(i,j)-\hat{A}(i_{1},j)+\varepsilon_{t}\geq 0)$ (58)
$=\mathbb{P}(\hat{A}(i,j)-A(i,j)-(\hat{A}(i_{1},j)-A(i_{1},j))\geq\Delta_{ij}^{\text{max}}-\varepsilon_{t})$ (59)
$\leq\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)}$ (60)
where we assume $A(i_{1},j)=\max_{i\in S_{x}}A(i,j)$ and $A(i,j_{1})=\min_{j\in S_{y}}A(i,j)$. To derive the inequality (58), we use the fact that the probability of the intersection of multiple events is always less than or equal to the probability of any of the individual events. Specifically, since we have $m$ actions in $S_{x}$, for any collection of events $e_{1},e_{2},\cdots,e_{m}$, we have $\mathbb{P}(e_{1}\cap e_{2}\cap\cdots\cap e_{m})\leq\mathbb{P}(e_{1})$, as in Lemma 3.
Then, we see that $\hat{A}(i,j)-A(i,j)-(\hat{A}(i_{1},j)-A(i_{1},j)$ is $\sqrt{2\sigma^{2}/k_{t}}$-subgaussian using subgaussian properties. Thus, we apply Lemma 4 to get inequality (60) since $\Delta_{ij}^{\text{max}}-\varepsilon_{t}\geq 0$ where $t$ is round $t_{ij}$ or later.
@@
$\mathbb{P}(E_{2})$ $=\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i,j)+\varepsilon_{t}\geq 0,\forall i^{\prime}\in S_{x})$ (61)
$\leq\mathbb{P}(\hat{A}(i,j_{1})-\hat{A}(i,j)+\varepsilon_{t}\geq 0)$ (62)
$=\mathbb{P}(\hat{A}(i,j_{1})-A(i,j_{1})-(\hat{A}(i,j)-A(i,j))\geq\Delta_{ij}^{\text{min}}-\varepsilon_{t})$ (63)
$\leq\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)}$ (64)
where $A(i,j_{1})=\min_{j\in S_{y}}A(i,j)$. We note that $\hat{A}(i,j_{1})-A(i,j_{1})-(\hat{A}(i,j)-A(i,j)$ is $\sqrt{2\sigma^{2}/k_{t}}$-subgaussian. Then, we apply Lemma 4, as we have $\Delta_{ij}^{\text{min}}-\varepsilon_{t}\geq 0$ for $t=t_{ij}$ or later rounds from the definition of $t_{ij}$.
Therefore, we write the probability of keeping an action pair $(i,j)$ as
$\mathbb{P}(E_{1}\cap E_{2})\leq\min\{\mathbb{P}(E_{1}),\mathbb{P}(E_{2})\}$ $\leq\min\Bigg\{\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)},\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)}\Bigg\}$ (65)
$=\exp{\Big(-\frac{k_{t}\max\{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2},(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}\}}{4\sigma^{2}}\Big)}$ (66)
$\leq\exp{\Big(-\frac{k_{t}((\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}+(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2})}{8\sigma^{2}}\Big)}$ (67)
$\leq\exp{\Big(-\frac{k_{t}(\Delta_{ij}-2\varepsilon_{t})^{2}}{16\sigma^{2}}\Big)}$ (68)
$<\exp{\Big(-\frac{k_{t}\varepsilon_{t}^{2}}{4\sigma^{2}}\Big)}$ (69)
$=\exp{\Bigg(-\frac{k_{t}\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{4\sigma^{2}}\Bigg)}$ (70)
$=\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}.$ (71)
For (67), we apply Lemma 1. Then , to write (68), we use Lemma 2. Using the definition of $t_{ij}$ and (54), $\varepsilon_{t}<\Delta_{ij}/4$ for rounds $t\geq t_{ij}$. Then, we derive the inequality (69). It is important to note that this probability bound holds only for rounds $t_{ij}$ or later. However, the probability of keeping a suboptimal action pair $(i,j)$ in round $t_{ij}$ also accounts for previous rounds, as it implies that the algorithm has not eliminated this pair in earlier rounds.
Therefore, the probability that a suboptimal action pair $(i,j)$ is not eliminated in round $t_{ij}$ or before is bounded by $\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}T}$. Then, the regret contribution is simply bounded by a worst case which is $T\Delta_{ij}^{*}$ for any suboptimal action pair $(i,j)\in S_{A_{1}}$.
@@
$k_{t_{ij}}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil\leq\Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil.$ (73)
@@
It leads to a misidentification of the NE, implying that an action pair $(i,j)$ should be better than the optimal one based on the current estimates. By the $\varepsilon$-NE property, we clearly have the following:
@@
However, it is not enough to keep a suboptimal action pair $(i,j)$. If the algorithm keeps a suboptimal action pair $(i,j)$ up to round $t_{*}$, the events specified in (56) must be satisfied with the parameter $\varepsilon_{t_{*}}$. Furthermore, we note that if the optimal action pair $(i^{*},j^{*})$ is eliminated in round $t_{*}$, it must fail to satisfy the $\varepsilon$-NE property as the following:
@@
We observe that all action pairs $(i,j)$ with $t_{ij}<t_{*}$ have already been eliminated in round $t_{ij}$ or earlier, which we already consider this condition in the previous case. Then, the optimal action pair can only be eliminated by a suboptimal action pair $(i,j)$ in round $t_{*}$ such that $t_{ij}\geq t_{*}$ since action pair $(i,j)$ must remain under consideration at the elimination round of the optimal action pair. This condition also implies that $\hat{\Delta}_{t_{ij}}\leq\hat{\Delta}_{t_{*}}$ because of the decreasing behavior of $\hat{\Delta}_{t}$ across rounds in the algorithm. Moreover, it is important to note that the algorithm keeps this action pair $(i,j)$ for at least one additional round following the elimination of the optimal pair as it is identified as a NE.
@@
We observe that the optimal action pair is eliminated by $(i,j)$ in some round $t_{*}$ where $t_{*}$ ranges from 0 to $t_{ij}$. Once this elimination occurs, the action pair action pair $(i,j)$ must remain active for at least one additional round, since it is selected as a misidentified NE. Hence, the algorithm necessarily keeps $(i,j)$ until at least round $t_{*}+1$. Furthermore, because $t_{*}\leq t_{ij}$, probability of keeping a suboptimal action pair $(i,j)$ in round $t_{*}+1$ can be bounded by $\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}+1}^{2}T}$. Thus, we can write the regret contribution of this case as
$\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}T\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}+1}^{2}T}$ $=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}T\frac{64\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}T}$ (77)
@@
where $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ and we apply (53) to get (78). We consider the action pairs in $S_{A_{2}}$ for this case, as this suboptimal action pair is identified as a NE. This is necessary because we need to account for regret from the commit phase in Algorithm 2 and the round $t_{ij}+1$ might be the round when the commit phase started.
@@
This implies that a suboptimal action pair is played during the commit phase of Algorithm 2 up to step $T$. Thus, we need to account for an additional regret contribution term given by a trivial upper bound $\max_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T$.